The length L in hours that a phone will work before it needs charge is normally distributed with a mean of 100 hours and a standard deviation of 15 hours.

i) Find p(L>127) and the 95th percentile of L

II)Alice is about to go for a 6hour journey. Given that it is 127 hours since last Alice charged her phone .find the probability that her phone will not need charging before the journey is completed.


1
Expert's answer
2021-08-19T07:42:03-0400

i) "L\sim N(\mu, \sigma^2)"

Given "\mu=100\ hr, \sigma=15\ hr"


"P(L>127)=1-P(L\leq 127)"

"=1-P(L\leq \dfrac{127-100}{15})"

"=1-P(Z\leq1.8)\approx0.035930"

"P(Z<z)=0.95"

"z\approx1.644854"

"\dfrac{x-100}{15}\approx1.644854"

"x\approx124.67"

ii)


"127+6=133"


"P(L>133)=1-P(L\leq 133)"

"=1-P(L\leq \dfrac{133-100}{15})"

"=1-P(Z\leq2.2)\approx0.013903"